# CAT-SAM: Conditional Tuning Network for Few-Shot Adaptation of Segmentation Anything Model

### v1

@@ Abstract
footnotetext: *indicates equal contribution.footnotetext: ${}^{\dagger}$ Corresponding author
The recent Segment Anything Model (SAM) has demonstrated remarkable zero-shot capability and flexible geometric prompting in general image segmentation. However, SAM often struggles when handling various unconventional images, such as aerial, medical, and non-RGB images. This paper presents CAT-SAM, a ConditionAl Tuning network that adapts SAM toward various unconventional target tasks with just few-shot target samples. CAT-SAM freezes the entire SAM and adapts its mask decoder and image encoder simultaneously with a small number of learnable parameters. The core design is a prompt bridge structure that enables decoder-conditioned joint tuning of the heavyweight image encoder and the lightweight mask decoder. The bridging maps the prompt token of the mask decoder to the image encoder, fostering synergic adaptation of the encoder and the decoder with mutual benefits. We develop two representative tuning strategies for the image encoder which leads to two CAT-SAM variants: one injecting learnable prompt tokens in the input space and the other inserting lightweight adapter networks. Extensive experiments over 11 unconventional tasks show that both CAT-SAM variants achieve superior target segmentation performance consistently even under the very challenging one-shot adaptation setup. Project page: https://xiaoaoran.github.io/projects/CAT-SAM
@@ 1 Introduction
Accurate segmentation is a fundamental computer vision task that plays a pivotal role in various applications such as robotics, autonomous driving, healthcare, earth observation, etc. The recently developed Segment Anything Model (SAM) [19], trained with 1.1 billion masks, has emerged as a momentous leap forward in image segmentation. By taking geometric prompts with points, boxes, or masks as input, SAM demonstrates remarkable zero-shot capability for general image segmentation, as well as great potential for fine-grained mask segmentation in different downstream tasks across a variety of contexts.
@@
On the other hand, SAM often struggles when dealing with various unconventional images [5, 14], such as challenging RGB images of aerial, medical, and intricate structural images, as well as non-RGB images like X-ray, Sonar, Synthetic Aperture Radar (SAR) images, etc., which greatly undermines its versatility and applicability as a foundational segmentation model while handling various downstream tasks. Several studies address this issue by fine-tuning SAM with a large number of target samples with mask annotations [17, 28]. However, such a tuning approach requires extensive masked images for each specialized field, posing a formidable task with poor scalability. How to effectively adapt SAM toward various unconventional images in a data-efficient manner has become an essential challenge in the segmentation field.
We exploit few-shot target samples for effective and efficient adaptation of SAM toward various downstream tasks. Specifically, we freeze SAM and expand its mask decoder lightly with a new prompt token and mask head (an MLP) as in [17], preserving SAM’s zero-shot capabilities and flexibility while achieving efficient adaptation concurrently. Such decoder-based tuning is well-suited for adapting normal natural images, benefiting from SAM’s well-trained image encoder on extensive natural images. However, it demonstrates degraded efficacy when dealing with unconventional images that often fall outside the scope of SAM’s training distribution, leading to suboptimal image segmentation.
We design CAT-SAM, a ConditionAl Tuning network for effective and data-efficient adaptation of SAM toward various unconventional image domains. CAT-SAM introduces a small number of trainable parameters to SAM’s image encoder and mask decoder for capturing domain-specific features and facilitating target segmentation. However, SAM’s image encoder is much heavier than its mask decoder, which directly leads to tuning imbalance and further suboptimal adaptation, especially under the few-shot setup with limited target samples. To this end, we design a prompt bridge structure that performs decoder-conditioned joint tuning by mapping the prompt tokens of the mask decoder to the image encoder, as depicted in Figure 1. With the communication between SAM’s image encoder and its mask decoder, CAT-SAM enables synergistic adaptation of the two network components which mitigates the tuning imbalance and improves few-shot SAM adaptation effectively.
The proposed prompt bridge can be seamlessly embedded into two representative tuning strategies: Prompt tuning approach  [25, 21, 16] that introduces learnable prompt tokens into the input space, and adapter-based approach [38, 39, 11, 27] that introduces lightweight adaptive networks. This directly leads to two CAT-SAM variants, namely, CAT-SAM-T and CAT-SAM-A, respectively. We extensively evaluate the two CAT-SAM variants over 11 unconventional datasets that SAM underperforms, ranging from challenging true-color images such as aerial, medical, and intricate structural images, to non-RGB images like X-ray, Sonar, and SAR images. Extensive evaluations show that both CAT-SAM variants achieve significant adaptation performance and consistently produce superior image segmentation even with very limited target samples.
@@
The major contributions of this work can be summarized in three aspects. First, we propose CAT-SAM, a conditional tuning network that enables effective and data-efficient adaptation of SAM toward various unconventional images. To this end, we design prompt bridge, a decoder-conditioned joint tuning structure that enables synergistic adaptation of the heavyweight image encoder and the lightweight mask decoder effectively. Second, we develop two CAT-SAM variants by embedding the prompt bridge into two representative tuning strategies, one introducing learnable prompt tokens in the input space and the other inserting lightweight adapter networks. Third, extensive experiments over 11 unconventional segmentation datasets show that CAT-SAM achieves superior image segmentation consistently even under the challenging one-shot setup.
@@ 2 Related Work
Figure 2: Overview of the proposed CAT-SAM. CAT-SAM keeps the whole SAM frozen while simultaneously tuning the image encoder and mask decoder for downstream adaptation. To address the tuning imbalance between these two network components, we introduce decoder-conditioned joint tuning through the design of Prompt Bridge structure, enabling synergetic and enhanced adaptation. We present two CAT-SAM variants: CAT-SAM-T in (a) and CAT-SAM-A in (b), achieved by integrating the prompt bridge with prompt-based and adapter-based tuning strategies for the image encoder, respectively. (c) illustrates two tailored prompt bridge structures, PB-T and PB-A.
Pretrained Foundation Models (PFMs) have gained significant popularity due to their exceptional zero-shot capability in addressing unseen data distributions and tasks, alongside their remarkable transferability in learned representations. PFMs have been extensively explored in Natural Language Processing (NLP), leading to the development of milestone models such as BERT [8], BART  [22], T5 [37], and the GPT series [34, 35, 2, 31]. Similarly, exploration of PFMs for visual-language has resulted in models such as CLIP [36], ALIGN [15], Flamingo [1], and Blip-2 [24]. Notably, SAM [19] was recently introduced as the first foundation model for visual segmentation, pre-trained on over 1.1 billion masks. This model allows users to interact by inputting various geometric prompts, demonstrating strong zero-shot capability across conventional image domains. Its potential across numerous applications represents a significant paradigm shift in both research and industry fields.
Parameter-Efficient Tuning for PFMs has become increasingly crucial due to the over-size model parameters and costly one-task-one-model deployment. One popular approach involves updating only the newly added learnable parameters while keeping the model backbone frozen. These lightweight parameters are often introduced in two ways: 1) as prompt tuning [53, 16, 18] by adding learnable tokens to input tokens at transformer layers, 2) as adapters [38, 39, 27, 10, 45] by integrating learnable lightweight sub-networks into each transformer layer. However, prior visual studies primarily focus on adapting single backbone modules [16, 27, 3] or dual encoder backbones for vision-language tasks [18, 52], which are suboptimal tuning solutions while applying to SAM due to its unique architecture of a heavyweight image encoder and a lightweight mask decoder. In contrast, we explore a novel conditional joint tuning strategy tailored to SAM, which facilitates its adaptation even with limited downstream data.
SAM for Downstream Tasks. The release of SAM has spurred several subsequent studies. Some concentrate on expanding semantic recognition capabilities [46, 43, 23, 20, 50], while others aim to create more lightweight SAM variants [51, 48] for faster computations. Several studies explore SAM’s adaptation in underperforming downstream scenarios, primarily by directly fine-tuning SAM’s mask decoder [4, 49, 7], which can severely degrade SAM’s zero-shot capabilities. The recent HQ-SAM [17] introduces a novel approach by freezing the entire pre-trained SAM and only prompt-tuning the mask decoder. It also introduces a new mask head, enabling to generation of the original SAM mask and adaptive mask simultaneously. Building upon HQ-SAM, our CAT-SAM proposes conditional joint tuning, facilitating synergistic tuning for both the image encoder and mask decoder, thereby enhancing adaptation in a data- and parameter-efficient manner.
@@ 3 Method
3.1 Preliminaries
Revisiting SAM. SAM [19] handles each image with three key modules: a heavyweight image encoder (i.e., ViT [9] as the backbone) that extracts image embeddings, a prompt encoder that encodes the geometric prompts (e.g., points, a box, or a coarse mask) to generate prompt embeddings, and a lightweight mask decoder that combines the two types of embeddings to predict segmentation masks. The released SAM model is trained on a super large-scale SA-1B dataset, which consists of over 11 million images as well as 1.1 billion automatically generated masks. SAM has demonstrated remarkable zero-shot capability while dealing with various conventional natural images, as well as superb flexibility in accepting various geometric prompt inputs. On the other hand, it often struggles while facing unconventional images, such as challenging RGB domains of aerial, medical, and intricate structural images, as well as non-RGB domains of X-ray, SAR, and Sonar images, which are often sparse or out-of-distribution falling outside of SAM’s training data. More details about SAM are presented in [19].
Revisiting HQ-SAM. HQ-SAM [17] keeps the entire pre-trained SAM frozen and introduces a learnable HQ-Output Token, which is concatenated with SAM’s output tokens and prompt tokens to form the input to SAM’s mask decoder. After forwarding through two decoder layers similarly as SAM’s output token, the updated HQ-Output Token is used for generating dynamic weights for a newly introduced three-layer MLP. Simultaneously, SAM’s mask decoder features are fused with the upsampled multi-scale features from its image encoder, leading to the HQ-feature. Finally, the HQ-feature and the three-layer MLP are combined through point-wise product to generate the target mask. On the other hand, HQ-SAM tunes SAM’s mask decoder only without adjusting its image encoder, leading to sub-optimal adaptation while handling various unconventional images that are sparse or out-of-distribution in SAM’s training data. Please refer to [17] for more details.
@@ 3.2 CAT-SAM
This subsection introduces CAT-SAM, a ConditionAl Tuning network that is designed for adapting SAM toward various downstream data. The objective is to adapt SAM efficiently with few-shot annotated target images only, meanwhile preserving SAM’s powerful zero-shot capability and geometric prompting flexibility.
Different from HQ-SAM that tunes the mask decoder only, the proposed CAM-SAM simultaneously tunes the image encoder to capture more target information from the few-shot target samples. However, the critical issue of the simultaneous tuning lies with a clear imbalance between the heavyweight image encoder (308.3 M parameters for ViT-L [9]) and the lightweight mask decoder (4.1 M). To tackle this challenge, we design a novel module called decoder-conditioned joint tuning that addresses the imbalance issue by building up a linkage between the encoding tuning and the decoding tuning. Specifically, we design a prompt bridge—a lightweight network that maps the HQ-Token of the mask decoder to the image encoder, leading to joint and balanced tuning of the image encoder and mask decoder as well as synergetic and more effective adaptation of SAM toward various downstream tasks.
We develop two CAT-SAM variants by embedding the proposed decoder-conditioned joint tuning into two prevalent tuning approaches. The first is CAT-SAM-T as shown in Figure 2 (a), which tunes the image encoder via prompt tuning. The second is CAT-SAM-A as shown in Figure 2 (b), which tunes the image encoder by introducing lightweight adapters. The decoder-conditioned joint tuning can be seamlessly embedded via two tailored Prompt Bridges that are denoted as PB-T and PB-A as illustrated in Figure 2 (c).
3.2.1 CAT-SAM-T
SAM’s Image encoder $\mathcal{V}$ with $K$ transformer layers $\{\mathcal{V}_{i}\}^{K}_{i=1}$, splits the input image $I$ of the size $(H,W)$ into $M$ fixed-size patches, and projects them into patch embeddings $E_{0}\in\mathbb{R}^{M\times d_{v}}$. Patch embeddings $E_{i}$ are input to the $(i+1)^{th}$ transformer layer $(\mathcal{V}_{i+1})$ and sequentially processed through $K$ transformer layers:
@@ 3.2.1 CAT-SAM-T
Prompt Tuning for the Image Encoder. As shown in Figure 2 (a), we introduce a set of learnable tokens $\{P_{i}\in\mathbb{R}^{d_{v}}\}^{b}_{i=1}$ in the image encoder alongside the input patch embeddings. These new learnable tokens are introduced in each transformer layer of the image encoder $\mathcal{V}$ and updated through the adaptation process:
@@
Decoder-Conditioned Joint Tuning. We employ a prompt bridge $\mathrm{PB}^{T}_{i}$, a three-layer MLP to project the HQ-Token $Q$ in the mask decoder to each transformer layer $i$ in the image encoder as the single learnable token replacing $P_{i}$, i.e., $\tilde{P}^{T}_{i}=\mathrm{PB}^{T}_{i}(Q)$. These mapped tokens are directly applied for prompt tuning the image encoder:
@@
The network structure of prompt bridge $\mathrm{PB}^{T}_{i}$ is depicted in Figure 2 (c).
3.2.2 CAT-SAM-A
Adapter Tuning for the Image Encoder. As illustrated in Figure 2 (b), we introduce Adapters to tune SAM’s image encoder in CAT-SAM-A. The adapter [11] is a lightweight sub-network that is inserted into each transformer layer, comprising a linear down-projection, a nonlinear activation function, a linear up-projection, and a residual connection.
High-frequency image information is also incorporated in CAT-SAM as the input of the adapter as in [27]. In this process, high-frequency components $I_{hfc}$ are first extracted from the input image $I$ using Fast Fourier Transform and its inverse. These components are then partitioned into small patches $I_{hfc}^{p}$ similar to the image patches. Subsequently, convolutional layers and linear layers are employed to independently project $I_{hfc}$ and the original image embeddings $E_{0}$ into $c$-dimensional features, yielding $F_{hfc}$ and $F_{pe}$, respectively. The result of element-wise addition of $F_{pe}$ and $F_{hfc}$ is then forwarded (as denoted ”EVP” in Figure 2 (b)) to the adapter $\mathrm{Adapt}_{i}$ for transformer layer $i$. The output is subsequently added element-wise to the original input of the transformer layer $i$:
$E_{i-1}=E_{i-1}+\mathrm{Adapt}_{i}(F_{pe}+F_{hfc}).$
Decoder-Conditioned Joint Tuning. To achieve decoder-conditioned joint tuning, we design a prompt bridge $\mathrm{PB}^{A}$ to map the HQ-Token $Q$ into an embedding $\tilde{P}^{A}=\mathrm{PB}^{A}(Q)$ of the same size as $F_{pe}$ and $F_{hfc}$. This embedding is element-wise added to the EVP before being fed to every adapter as follows:
@@ 3.2.2 CAT-SAM-A
As illustrated in Figure 2 (c), PB-A employs $c$ different linear down-projection layers to generate embeddings independently. After GELU layers, these embeddings are up-projected to the dimension of $1\times M$ ($M$ is the spatial size of the patch) with a shared linear layer to mitigate the computation cost. Finally, the up-projected embeddings are concatenated to produce $\tilde{P}^{A}$.
Table 1 compares the amount of trainable parameters between the SAM (based on the ViT-L [9]) and our CAT-SAM that is built upon the SAM. It can be observed that both CAT-SAM variants introduce very limited additional parameters, yet yield significant performance improvements in various downstream segmentation tasks as to be detailed in Section 4.
Model Train. Params. (M) Percent.
@@
Table 1: Trainable parameter comparison between ViT-L [9] based SAM and CAT-SAM variants.

### v3

# CAT-SAM: Conditional Tuning for Few-Shot Adaptation of Segment Anything Model

@@ Abstract
[…]
The Segment Anything Model (SAM) has demonstrated remarkable zero-shot capability and flexible geometric prompting in general image segmentation. However, it often struggles in domains that are either sparsely represented or lie outside its training distribution, such as aerial, medical, and non-RGB images. Recent efforts have predominantly focused on adapting SAM to these domains using fully supervised methods, which necessitate large amounts of annotated training data and pose practical challenges in data collection. This paper presents CAT-SAM, a ConditionAl Tuning network that explores few-shot adaptation of SAM toward various challenging downstream domains in a data-efficient manner. The core design is a prompt bridge structure that enables decoder-conditioned joint tuning of the heavyweight image encoder and the lightweight mask decoder. The bridging maps the domain-specific features of the mask decoder to the image encoder, fostering synergic adaptation of both components with mutual benefits with few-shot target samples only, ultimately leading to superior segmentation in various downstream tasks. We develop two CAT-SAM variants that adopt two tuning strategies for the image encoder: one injecting learnable prompt tokens in the input space and the other inserting lightweight adapter networks. Extensive experiments over 11 downstream tasks show that CAT-SAM achieves superior segmentation consistently even under the very challenging one-shot adaptation setup.
Keywords:
Segment Anything Model Few-shot Learning Parameter Efficient Tuning
@@ 1 Introduction
Accurate image segmentation is a fundamental computer vision task that plays a pivotal role in various applications such as robotics, autonomous driving, healthcare, earth observation, etc. The recently developed Segment Anything Model (SAM) [23], trained with 1.1 billion masks, has emerged as a momentous leap forward in image segmentation. By taking geometric prompts with points, boxes, or masks as input, SAM demonstrates remarkable zero-shot capability for general image segmentation, as well as great potential for fine-grained mask segmentation in different downstream tasks across a variety of contexts.
@@
On the other hand, SAM often fails to generate high-quality predictions when dealing with domains that are either sparsely represented or lie outside its training distribution [5, 19], such as challenging RGB images including aerial, medical, and intricate structural images, as well as non-RGB images like X-ray, Sonar, Synthetic Aperture Radar (SAR) images, etc. This greatly undermines the versatility and applicability of SAM as a foundational segmentation model while handling various downstream tasks. Several studies attempted to address this issue by fine-tuning SAM with a large number of target samples with mask annotations [21, 30]. However, they require extensive masked images for each specialized field, posing a formidable task with poor scalability. How to effectively adapt SAM toward various downstream tasks in a data-efficient manner has become an essential challenge in the segmentation field.
We exploit few-shot target samples for effective and efficient adaptation of SAM toward various downstream tasks. Inspired by the recent advancements in parameter-efficient tuning for foundation models [1, 13], we freeze the entire SAM and lightly expand its image encoder and mask decoder with a tiny amount of learnable parameters. This expansion preserves SAM’s zero-shot capabilities and flexibility while capturing domain-specific features that are essential for the segmentation of downstream data. However, SAM’s image encoder is much larger than its mask decoder, which directly leads to tuning imbalance and further suboptimal adaptation, especially under the limited supervision of few-shot target samples.
To address this challenge, we design CAT-SAM, a novel ConditionAl Tuning network for effective and data-efficient adaptation of SAM. The key idea is to exploit the lightweight decoder tuning to guide the heavyweight encoder tuning and formulate decoder-conditioned joint tuning of both as depicted in  Fig. 1. To this end, we design a prompt bridge structure that maps the domain-specific features from the mask decoder to the image encoder. With the collaboration between the two partners, CAT-SAM enables synergistic adaptation of both which mitigates the tuning imbalance and improves the few-shot SAM adaptation effectively. In addition, the prompt bridge can be seamlessly embedded into two representative tuning approaches, including prompt tuning [27, 25, 20] that introduces learnable prompt tokens into the input space and adapters [37, 38, 15, 29] that leverage lightweight adaptive networks. This directly leads to two CAT-SAM variants, namely, CAT-SAM-T and CAT-SAM-A. Extensive evaluations over 11 diverse and challenging downstream tasks show that both variants achieve superior adaptation and segmentation with few-shot target samples only.
@@
The major contributions of this work can be summarized in three aspects. First, we propose CAT-SAM, a conditional tuning network that enables effective and data-efficient adaptation of SAM toward various challenging downstream domains. We design prompt bridge within CAT-SAM, a decoder-conditioned joint tuning structure that enables synergistic and data-efficient adaptation of the heavyweight image encoder and the lightweight mask decoder effectively. Second, we develop two CAT-SAM variants by embedding the prompt bridge into two representative tuning strategies, one introducing learnable prompt tokens in the input space and the other inserting lightweight adapter networks. Third, extensive experiments over 11 diverse segmentation datasets show that CAT-SAM achieves superior image segmentation consistently even under the challenging one-shot setup.
@@ 2 Related Work
Parameter-Efficient Tuning for foundation models has become increasingly crucial due to the over-size model parameters and costly one-task-one-model deployment. One popular approach involves updating only the newly added learnable parameters while keeping the model backbone frozen. These lightweight parameters are often introduced in two ways: 1) as prompt tuning [58, 20, 22, 12] by adding learnable tokens to input tokens at transformer layers, 2) as adapters [37, 38, 29, 11, 49] by integrating learnable lightweight sub-networks into transformer layers. However, prior visual studies primarily focus on adapting single backbone modules [20, 29, 3] or dual encoder backbones for vision-language tasks [22, 57], which are suboptimal tuning solutions while applying to SAM due to its particular architecture of a heavyweight image encoder and a lightweight mask decoder. In contrast, we explore a novel conditional joint tuning strategy tailored to SAM, which facilitates its adaptation even with limited downstream data.
SAM for Downstream Tasks. The release of SAM has spurred several subsequent studies. Some concentrate on expanding semantic recognition capabilities [50, 43, 26, 24, 54], while others aim to create more lightweight SAM variants [55, 52, 48] for faster computations. Several studies explore SAM’s adaptation in underperforming downstream scenarios, primarily by directly fine-tuning SAM’s mask decoder [4, 53, 8, 46, 41, 30], which can severely degrade SAM’s zero-shot capabilities. The recent HQ-SAM [21] freezes the entire pre-trained SAM and prompt-tunes the mask decoder with a new mask head for adaptive mask prediction. However, these adaptation studies rely on large target datasets with a considerable number of annotated images, which are often costly to collect. In contrast, our CAT-SAM achieves superior adaptation with only few-shot target samples, significantly reducing adaptation costs and expanding the applicability of SAM as a visual foundation model.
@@ 3 Method
3.1 Preliminaries of SAM
SAM [23] handles each image with three key modules: a heavyweight image encoder (i.e., ViT [9] as the backbone) that extracts image embeddings, a prompt encoder that encodes the geometric prompts (i.e., points, a box, or a coarse mask) to generate prompt embeddings, and a lightweight mask decoder that combines the two types of embeddings to predict segmentation masks. The released SAM model is trained on a super large-scale SA-1B dataset, which consists of over 11 million images as well as 1.1 billion automatically generated masks. SAM has demonstrated remarkable zero-shot capability while dealing with various conventional natural images, as well as superb flexibility in accepting various geometric prompt inputs. On the other hand, it often struggles while facing challenging downstream domains that are often sparse or falling outside of its training data distribution, such as RGB domains of aerial, medical, and intricate structural images, as well as non-RGB domains of X-ray, SAR, and Sonar images. More details about SAM are presented in [23].
@@ 3.2 CAT-SAM
This subsection introduces CAT-SAM, a ConditionAl Tuning network that is designed for adapting SAM toward various under-performed downstream data. The objective is to adapt SAM efficiently with only few-shot annotated target images, meanwhile preserving SAM’s powerful zero-shot capability and geometric prompting flexibility.
CAT-SAM freezes the entire SAM and facilitates the simultaneous tuning of SAM’s image encoder and mask decoder to capture representative information from few-shot target samples for effective adaptation of SAM. However, a critical issue with simultaneous tuning arises from the clear imbalance between the heavyweight image encoder (308.3 M parameters for ViT-L [9]) and the lightweight mask decoder (4.1 M), which often leads to sub-optimal adaptation while only few-shot target samples are available. To tackle this challenge, we design decoder-conditioned joint tuning, a novel tuning strategy that mitigates the imbalance by establishing a linkage between the encoding tuning and the decoding tuning. Specifically, we design prompt bridge—a lightweight network that maps the domain-specific features from the mask decoder to the image encoder. The bridge design facilitates joint and balanced tuning of both network components, leading to synergistic and effective adaptation of SAM with few-shot and even a single target sample.
We develop two CAT-SAM variants by integrating the proposed decoder-conditioned joint tuning into two prevalent tuning approaches, namely CAT-SAM-T and CAT-SAM-A, as illustrated in Fig. 2.
Figure 2: Overview of CAT-SAM. CAT-SAM keeps the whole SAM frozen while simultaneously tuning the image encoder and mask decoder for downstream adaptation. To address the tuning imbalance between these two network components, we introduce decoder-conditioned joint tuning through the design of Prompt Bridge structure, enabling synergetic and enhanced adaptation. We present two CAT-SAM variants: CAT-SAM-T in (a) and CAT-SAM-A in (b), achieved by integrating the prompt bridge with prompt-based and adapter-based tuning strategies for the image encoder, respectively. (c) illustrates two tailored prompt bridge structures, PB-T and PB-A.
3.2.1 Tuning Image Encoder.
We first brief SAM’s image encoder and then introduce two tuning approaches in our CAT-SAM, including prompt tuning in CAT-SAM-T and lightweight adapters in CAT-SAM-A.
  • $\bullet$
    SAM’s Image Encoder, denoted as $\mathcal{V}$ with $K$ transformer layers $\{\mathcal{V}_{i}\}^{K}_{i=1}$, first splits the input image $I$ of the size $(H,W)$ into $M$ fixed-size patches, and then projects them into patch embeddings $E_{0}\in\mathbb{R}^{M\times d_{v}}$. Patch embeddings $E_{i-1}$ are input to the $i^{th}$ transformer layer $(\mathcal{V}_{i})$ and sequentially processed through $K$ transformer layers:
@@
  • $\bullet$
    CAT-SAM-T: As illustrated in Fig. 2 (a), we introduce a set of learnable tokens $\{P_{i}\in\mathbb{R}^{d_{v}}\}^{K-1}_{i=0}$ in the image encoder alongside the input patch embeddings. These new learnable tokens are introduced in each transformer layer of the image encoder $\mathcal{V}$ and updated through the adaptation process:
@@
  • $\bullet$
    CAT-SAM-A: As illustrated in Fig. 2 (b), we introduce Adapters [15] to tune SAM’s image encoder in CAT-SAM-A. The adapter is a lightweight sub-network that is inserted into each transformer layer, comprising a linear down-projection, a nonlinear activation function, a linear up-projection, and a residual connection. High-frequency image information is also incorporated in CAT-SAM as the input of the adapters as in [29]. In this process, high-frequency components $I_{hfc}$ are first extracted from the input image $I$ and then reversed to the space domain using the Fast Fourier Transform and its inverse. These components are then partitioned into small patches $I_{hfc}^{p}$ similar to the image patches. Subsequently, convolutional layers and linear layers project $I_{hfc}$ and the original image embeddings $E_{0}$ into $c$-dimensional features independently, yielding $F_{hfc}$ and $F_{pe}$, respectively. The result of element-wise addition of $F_{pe}$ and $F_{hfc}$ is then forwarded (as denoted "EVP" in Fig. 2 (b)) to the adapter $\mathrm{Adapt}_{i}$ for transformer layer $i$. Finally, the output is added element-wise to the original input of the transformer layer $i$:
    $E_{i-1}=E_{i-1}+\mathrm{Adapt}_{i}(F_{pe}+F_{hfc}).$
3.2.2 Tuning Mask Decoder.
We tune the mask decoder of SAM similarly as [21]. Specifically, we freeze the entire pre-trained decoder of SAM and introduce a learnable CAT-Token that is concatenated with SAM’s original output tokens and prompt tokens to form the input to SAM’s mask decoder. The CAT-Token is updated during training but remains fixed during inference, the same as SAM’s output tokens. After passing through two decoder layers, the updated CAT-Token is used to generate dynamic weights for a newly introduced three-layer MLP. Simultaneously, SAM’s mask decoder features are fused with the image features decoded by the image encoder. Finally, the fused features and the three-layer MLP are combined through the dot product to generate the target mask. Please refer to the appendix for more details about the decoder structures in CAT-SAM.
3.2.3 Decoder-Conditioned Joint Tuning.
The proposed decoder-conditioned joint tuning can be seamlessly embedded via two implementations of the proposed prompt bridging ideas, i.e., PB-T and PB-A as illustrated in Fig. 2 (c).
  • $\bullet$
    CAT-SAM-T: The prompt bridge $\mathrm{PB}^{T}_{i}$ is a two-layer MLP. It projects the CAT-Token $Q$ in the mask decoder to each transformer layer $i$ in the image encoder and outputs a single learnable token replacing $P_{i}$ by $\tilde{P}^{T}_{i}=\mathrm{PB}^{T}_{i}(Q)$. These mapped tokens are directly applied for prompt tuning of the image encoder as follows:
@@
  • $\bullet$
    CAT-SAM-A: The prompt bridge $\mathrm{PB}^{A}$ maps the CAT-Token $Q$ into an embedding $\tilde{P}^{A}=\mathrm{PB}^{A}(Q)$ of the same size as $F_{pe}$ and $F_{hfc}$. $\tilde{P}^{A}$ is then element-wise added to the EVP before being fed to every adapter as follows:
@@
    As illustrated in Fig. 2 (c), PB-A employs $c$ different linear down-projection layers to generate embeddings independently. After GELU layers, these embeddings are up-projected to the dimension of $1\times M$ ($M$ is the spatial size of the patch) with a shared linear layer to mitigate the computation cost. Finally, the up-projected embeddings are concatenated to produce $\tilde{P}^{A}$.
Table 1: Comparisons of trainable parameters between ViT-L [9] based SAM and two variants of the proposed CAT-SAM.
Model Trainable Parameters (M) Trainable Params vs. SAM
@@
Tab. 1 compares the amount of trainable parameters between the SAM (implemented with the ViT-L [9]) and our CAT-SAM that is built upon the SAM. It can be observed that both CAT-SAM variants introduce very limited additional parameters, yet yield significant performance improvements in various downstream segmentation tasks as to be detailed in Section 4.